Hi I am having trouble with the following problem.
Determine whether the following two planes 4x+y-z=3 and 2x-5y+3z=2 are parallel, orthogonal, coincident (that is, the same) or none of these?
Help with this would be much appreciated.
Thanks, Judy

Answer by Alan3354(69443)

Determine whether the following two planes 4x+y-z=3 and 2x-5y+3z=2 are parallel, orthogonal, coincident (that is, the same) or none of these?
------------
Find the vectors normal to each plane:
4x+y-z=3 --> 4i + j - k = normal vector A
2x-5y+3z=2 --> 2i - 5j + 3k = normal vector B
-----------
The dot product of the vectors is |A|*|B|cos(a) where a is the angle between the planes.
The magnitude of A, |A|, is sqrt(16 + 1 + 1) = 2sqrt(3)
The magnitude of B, |B|, is sqrt(4 + 25 + 9) = sqrt(38)
-----------
Dot product = 4*2 + 1*(-5) -1*3 = 0
Since neither A or B is 0, the cos(a) = 0
a = 90º, they're orthogonal